# Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding Zusammenfassung

### v1

@@
Area Selective Chemical Vapor Deposition of Gold by Electron Beam Seeding
Zusammenfassung
@@
Introduction
@@
Gold is a key metal in microelectronics due to its low resistivity, resistance to electromigration and chemical inertness. These properties make it an excellent choice for stable and reliable electrical contacts even under challenging environmental conditions 1. Chemical vapour deposition (CVD) is a technologically significant option for fabricating gold contacts or other metallic functional structures. CVD utilises the thermal decomposition of precursor molecules in a gaseous state 2. Molecules reaching the surface are thermally decomposed, leaving the central metal atom on the surface. If the decomposition is favored kinetically or thermodynamically on certain surfaces compared to others, the ensuing variation in reaction rates can be exploited for area selective deposition 2. This has been shown for various metals, e.g. Au on W 3, Au on Cr 4, or Cu on Pt 5. All of the aforementioned approaches require sophisticated lithography methods to define the growth surfaces, which can restrict the selection of substrate materials, ultimately constraining the degrees of freedom in device design 3, 6. Consequently, an area selective deposition technique that operates without the need for lithography, and is independent of the substrate type, would be highly advantageous. One option is to provide a localized energy input for CVD, which is frequently accomplished through laser irradiation 7, 8. However, the spatial resolution achievable through this method is typically restricted by the wavelength of the laser light employed. Although there has been recent evidence that this issue can be addressed in thermally decoupled 3D gold structures 9, the range of achievable structure geometries remains restricted. This is also applicable to local energy input through resistive heating 10. An alternative approach uses beams of charged particles, often electrons 11, 12. Instead of supplying additional energy during the CVD process, electrons can also be used to activate a specific surface 13, 14 or to deposit a seed layer for preferential decomposition of precursor molecules either by CVD 12, 15 or by atomic layer deposition (ALD) 16, 17. Seeding using a focused electron beam possesses significant potential, as direct electron beam writing is capable of operating on any surface topography that is even weakly conductive 18. Additionally, it provides the possibility for unparalleled spatial resolution 19. If the seeds previously deposited exhibit autocatalytic activity 12, 15, 20, 21, the temperature necessary for decomposition decreases considerably. This technique efficiently limits the CVD metal growth to a specific area, while also reducing the thermal load on the substrate. However, there is currently no proof of electron beam seeding resulting in autocatalytic growth for gold.
Results and Discussion
The principle of the proposed process is depicted in Figure 1. A seed layer is deposited by the impact of a focused electron beam on a sample surface at room or elevated temperature. The electron beam is then switched off while the molecular gas flow is maintained for a period of time. Subsequent investigations show autocatalytic growth of pure gold at both seeding temperatures.Electron Beam Seeding
Autocatalytic Growth of Gold
In order to achieve autocatalytic activity of the seed layer, two conditions must be met. Firstly, the previously deposited metallic nuclei must have a clean surface to participate in the reaction. Therefore, ultra-high vacuum (UHV) conditions are commonly used for this type of experiment. Additionally, all final reaction products that result from ligand cleavage must be volatile enough to desorb intactly 31. Secondly, the growth kinetics at the nucleus must promote chemisorption of the precursor molecules through the removal of the initial ligand with activation energy lower than the standard bond dissociation energies 15. This condition applies to some carbonyls, e.g. area selective CVD by electron beam seeding has been demonstrated for the deposition of transition metals using precursors like Fe(CO) 12, 21, Co(CO)NO 20, and Cr(CO) 15. In some cases high vacuum (HV) conditions were sufficient to observe autocatalysis, e.g. for Fe(CO) 40, but in others clean metal surfaces were additionally required excluding electron beam writing for seeding, e.g. for Co(CO) 41. Under ultra high vacuum conditions, NH dosing was found to inhibit autocatalytic growth, which could subsequently be reactivated by electron beam seeding, providing an additional control option in area selective CVD utilising Fe(CO) 21. Generally, carbonyls possess the thermodynamically stable CO ligand, making them the preferable choice for autocatalytically enhanced growth processes. In contrast, organic ligands such as the frequently used -diketonates (here acac) are not expected to have such favorable desorption properties as mentioned above. The autocatalytic growth experiments were performed in the same vacuum chamber. After depositing a square array for one temperature, the electron beam was switched off and the GIS nozzle was opened again for 30 min to deliver precursor molecules. Scanning electron (SE) micrographs for investigation were taken only after cooling the stage to room temperature and recovering the background pressure to the base pressure of about mbar. The experimental matrix continued with the next array, covering temperatures ranging from 30 °C up to 90 °C at intervals of 10 K. The growth results are shown in Figure 2(b) where carbon appears as a dark contrast and gold as a bright contrast, which is a typical characteristic of SE imaging. At 60 °C, no observable difference was noted when compared to the seed arrays deposited initially. However, at 70 °C, initial signs of CVD growth became evident. The contrast shifted from dark (carbon-rich) to bright (gold-rich), becoming more pronounced as the growth temperature increased. Squares with a greater number of repeats in the electron beam seeding are brighter, corresponding to more pronounced gold growth. This can be attributed to an increased density of gold nuclei available with deposition time. The optimal balance between growth rate and selectivity is achieved at a temperature of 80 °C. Raising the temperature further enhances the growth rate, but decreases selectivity due to the initiation of growth in the halo regions around the actual patterned squares. This phenomenon is demonstrated in Figure 2(c), a high-resolution SE micrograph of the square seeded at 90 °C with 20 repetitions at 5 nm pitch. Although there are no primary electrons in the halo, there are still secondary and backscattered electrons present with a significantly lower flux than in the deposition region 38. Consequently, numerous repetitions in seeding at 80 °C can also diminish the precision of square edges. Interestingly, electron beam seeding also works at room temperature. However, as anticipated, the higher desorption rate of ligands during seeding at elevated stage temperature suppresses carbon co-deposition, resulting in higher densities of catalytically active seeds. This is illustrated in Figures 1(a) and (b). Therefore, the films grown from seed layers deposited at room temperature show a significant carbon contrast, while at the elevated stage temperature the seed density is already high enough to form a quasi-continuous gold film, cf. Figure 2(b) middle row. For the tested pitch values and numbers of repeats, the growth of the autocatalytic film at a certain temperature is only influenced by the total electron dose. When comparing SE contrast and topography, it is evident that the seeds, which were deposited with different pitches and repeats but the same total dose, display the same thickness and SE contrast (grain size) after growth. The optimum dose is reached at about 140 pC/µm or 875 electrons per nm. Additionally, the process exhibits exceptional selectivity as gold growth is restricted to the presence of gold nuclei under these temperatures. It is worth noting, however, that a substantial decrease in CVD growth rate was observed over time, resulting in saturation for attainable thicknesses at approximately 6.5 nm. This could be due to surface poisoning, i.e. carbon coverage of the active gold nuclei due to residual hydrocarbons in the vacuum chamber and/or low desorption rates of the ligand fragments. In this instance, an improvement in the background vacuum would be promising. Another potential reason is the nuclei changing in size during autocatalytic growth. Size effects can have a significant impact on the catalytic activity of metallic nanoparticles, especially in the case of gold 42. It is worth mentioning that the observed decreasing CVD growth rate is not an obstacle to technological implementation, as seeding and growth can be easily iterated several times (as shown in Figure SI.3).
Beam-induced Heating and Autocatalysis
One of the main benefits of using a tightly focused electron beam to decompose molecules from the gas phase is its extreme 3D flexibility, resolution and shape fidelity 32. Generally, FEBID is considered to be stimulated by secondary electrons which dissociate the physisorbed precursor molecules. Electron-induced chemistry involves various reaction pathways, which depend on the local availability of electrons, molecules, and co-reactants 25, 31. Beam-induced heating is also a significant factor. As the growth surface locally heats up, the dynamics of the supplied precursor molecules and fragments will shift, e.g. towards faster diffusion and higher desorption rates. Estimates show that even slight changes in surface temperature, around 10 K, can significantly impact the vertical growth rate during 3D FEBID 43, 44, 45, 46, which typically operates in the reaction-rate limited growth regime. However, there is no experimental data available regarding the actual local temperature changes that may occur. Here, we exploit the onset of autocatalytic CVD during the growth of one-dimensional vertical nanostructures (pillars) under steady-state growth conditions as a means of in situ temperature measurement.
Conclusions
In this work, area selective CVD of gold by electron beam seeding is demonstrated for the first time. Gold nuclei, acting as a seed layer, are deposited through a maskless procedure by means of a focused electron beam, followed by gentle heating of the substrate, which triggers autocatalytic gold growth only in the seeded regions. The use of a focused electron beam ensures unparalleled spatial resolution, unhampered by mask usage typical in electron beam lithography and rendering complex multi-stage lithographic processing unnecessary. Seeding and subsequent growth both use the same precursor compound, Au(acac)Me, which is commercially available, long-term stable and non-toxic. The technique has been demonstrated to operate at low temperatures of approximately 80 °C, more than 200 K below the typical substrate temperature used in the established CVD process for this precursor. Moreover, it demands solely a high vacuum, making technological implementation remarkably straightforward. Any commercially available scanning electron microscope can be used for this method if it is equipped with a gas injection system and a heating stage. In the future, this process could be seamlessly integrated into conventional CVD reactors by equipping them with an electron gun featuring programmable position control. Unlike many other area selective CVD processes, direct electron beam writing for seeding is not affected by the surface material or topography. To demonstrate this capability, the same precursor was used to fabricate vertical nanostructures with high aspect ratios using the focused electron beam. In addition to achieving autocatalytically active seeds in 3D, the first quantitative measurement of the temperature change during direct electron beam writing of 3D structures was performed. This quantitative determination of the up to now inaccessible local temperature evolution is a major breakthrough as it enables improved modelling for accurate shape 3D nanoprinting through FEBID within the additive manufacturing field at a sub-100 nm scale 47, 49, 44, 45, 46. To summarise, this work presents a new, technologically powerful and easy to implement approach to area selective gold deposition. The low temperature necessary makes it highly suitable for processing innovative flexible devices that rely on heat-sensitive polymers, including flexible electronics 50 and flexible solar cells 51. Another exciting area of application is utilizing the maskless direct write seeding on 3D architectures, which enables the fabrication of nanostructures or localized electric contacts on intricate 3D devices.Acknowledgements
This work was funded by the German Research Association under grant agreement no. HO 5461/3-1 and by the Swiss National Fund by the COST-SNF project IZCOZ0_205450. We want to furthermore acknowledge support by the EU COST action CA 19140 ’FIT4NANO’ (www-fit4nano.eu). The authors are particularly grateful to Caroline Hain for her support with the English language.Conflict of Interest
Authors declare no conflict of interest. Keywords:Literatur
- [1] T. H. Baum, P. B. Comita, Chemical Vapor Deposition of Gold and Silver, chapter 6, pages 303–327, John Wiley & Sons, Ltd 1994.
- [2] T. T. Kodas, M. J. Hampden-Smith, Overview of Metal CVD, chapter 9, pages 429–498, John Wiley & Sons, Ltd 1994.
- [3] S. O. Colgate, G. J. Palenik, V. E. House, D. W. Schoenfeld, Journal of Vacuum Science & Technology A 1990, 8, 1411.
- [4] M. M. Banaszak Holl, P. F. Seidler, S. P. Kowalczyk, F. R. McFeely, Applied Physics Letters 1993, 62, 1475.
- [5] B. Lecohier, B. Calpini, J. Philippoz, H. van den Bergh, Journal of Applied Physics 1992, 72, 2022.
- [6] W. L. Gladfelter, Chemistry of Materials 1993, 5, 1372.
- [7] T. H. Baum, C. R. Jones, Applied Physics Letters 1985, 47, 538.
- [8] S. J. C. Irvine, D. Lamb, Photo-assisted CVD, in Chemical Vapour Deposition: Precursors, Processes and Applications, The Royal Society of Chemistry 2008.
- [9] J. Lasseter, P. D. Rack, S. J. Randolph, Nanomaterials 2023, 13.
- [10] F. Porrati, S. Barth, G. C. Gazzadi, S. Frabboni, O. M. Volkov, D. Makarov, M. Huth, ACS Nano 2023, 17, 4704, pMID: 36826847.
- [11] R. R. Kunz, T. E. Allen, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1987, 5, 1427.
- [12] R. R. Kunz, T. M. Mayer, Applied Physics Letters 1987, 50, 962.
- [13] M.-M. Walz, M. Schirmer, F. Vollnhals, T. Lukasczyk, H.-P. Steinrück, H. Marbach, Angewandte Chemie International Edition 2010, 49, 4669.
- [14] K. Muthukumar, H. O. Jeschke, R. Valentí, E. Begun, J. Schwenk, F. Porrati, M. Huth, Beilstein Journal of Nanotechnology 2012, 3, 546.
- [15] R. R. Kunz, T. M. Mayer, Journal of Vacuum Science & Technology B: Microelectronics Processing and Phenomena 1988, 6, 1557.
- [16] A. J. M. Mackus, J. J. L. Mulders, M. C. M. van de Sanden, W. M. M. Kessels, Journal of Applied Physics 2010, 107, 116102.
- [17] A. Mameli, B. Karasulu, M. A. Verheijen, B. Barcones, B. Macco, A. J. M. Mackus, W. M. M. E. Kessels, F. Roozeboom, Chemistry of Materials 2019, 31, 1250.
- [18] I. Utke, P. Hoffmann, J. Melngailis, J. Vac. Sci. Technol. B 2008, 26, 1197.
- [19] W. F. van Dorp, C. W. Hagen, P. A. Crozier, P. Kruit, Nanotechnology 2008, 19, 225305.
- [20] F. Vollnhals, M. Drost, F. Tu, E. Carrasco, A. Späth, R. H. Fink, H.-P. Steinrück, H. Marbach, Beilstein Journal of Nanotechnology 2014, 5, 1175.
- [21] P. Martinović, L. Barnewitz, M. Rohdenburg, P. Swiderek, Journal of Vacuum Science & Technology A 2023, 41, 033207.
- [22] E. Feurer, H. Suhr, Applied Physics A 1987, 44, 171.
- [23] C. E. Larson, T. H. Baum, R. L. Jackson, MRS Online Proceedings Library (OPL) 1986, 75, 721.
- [24] A. Botman, J. J. L. Mulders, C. W. Hagen, Nanotechnology 2009, 20, 17pp.
- [25] S. Barth, M. Huth, F. Jungwirth, J. Mater. Chem. C 2020, 8, 15884.
- [26] J. J. L. Mulders, J. M. Veerhoek, E. G. T. Bosch, P. H. F. Trompenaars, Journal of Physics D: Applied Physics 2012, 45, 475301.
- [27] I. Utke, P. Hoffmann, B. Dwir, K. Leifer, E. Kapon, P. Doppelt, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures 2000, 18, 3168.
- [28] A. Marashdeh, T. Tiesma, N. J. C. van Velzen, S. Harder, R. W. A. Havenith, J. T. M. De Hosson, W. F. van Dorp, Beilstein Journal of Nanotechnology 2017, 8, 2753.
- [29] D. McIntosh, G. A. Ozin, Inorganic Chemistry 1977, 16, 51.
- [30] L. Jiang, Q. Xu, The Journal of Physical Chemistry A 2005, 109, 1026.
- [31] I. Utke, P. Swiderek, K. Höflich, K. Madajska, J. Jurczyk, P. Martinović, I. Szymańska, Coordination Chemistry Reviews 2022, 458, 213851.
- [32] R. Winkler, J. D. Fowlkes, P. D. Rack, H. Plank, Journal of Applied Physics 2019, 125, 210901.
- [33] O. Sqalli, I. Utke, P. Hoffmann, F. Marquis-Weible, Journal of Applied Physics 2002, 92, 1078.
- [34] J. J. L. Mulders, L. M. Belova, A. Riazanova, Nanotechnology 2010, 22, 055302.
- [35] K. Höflich, R. B. Yang, A. Berger, G. Leuchs, S. Christiansen, Adv. Mater. 2011, 23, 2657.
- [36] R. Winkler, F. P. Schmidt, U. Haselmann, J. D. Fowlkes, B. B. Lewis, G. Kothleitner, P. D. Rack, H. Plank, ACS Applied Materials and Interfaces 2017, 9, 8233.
- [37] D. Kuhness, A. Gruber, R. Winkler, J. Sattelkow, H. Fitzek, I. Letofsky-Papst, G. Kothleitner, H. Plank, ACS Applied Materials and Interfaces 2021, 13, 1178.
- [38] K. Höflich, J. Jurczyk, M. Puydinger, Y. Zhang, C. Kapusta, I. Utke, M. Puydinger dos Santos, M. Götz, C. Guerra-Nuñez, J. Best, C. Kapusta, I. Utke, ACS Appl. Mater. Interfaces 2017, 9, 24071.
- [39] P. Martinović, M. Rohdenburg, A. Butrymowicz, S. Sarigül, P. Huth, R. Denecke, I. B. Szymanska, P. Swiderek, Nanomaterials 2022.
- [40] G. Hochleitner, H. D. Wanzenboeck, E. Bertagnolli, Journal of Vacuum Science & Technology B: Microelectronics and Nanometer Structures Processing, Measurement, and Phenomena 2008, 26, 939.
- [41] R. Córdoba, J. Sesé, M. Ibarra, J. De Teresa, Applied Surface Science 2012, 263, 242.
- [42] M. Haruta, Catal. Today 1997, 36, 153 .
- [43] S. J. Randolph, J. D. Fowlkes, P. D. Rack, Journal of Applied Physics 2005, 97, 124312.
- [44] E. Mutunga, R. Winkler, P. D. Rack, H. Plank, J. D. Fowlkes, ACS Nano 2019, 13, 5198 .
- [45] L. Skoric, D. Sanz-Hernández, F. Meng, C. Donnelly, S. Merino-Aceituno, A. Fernández-Pacheco, Nano Letters 2020, 20, 184.
- [46] J. D. Fowlkes, R. Winkler, P. D. Rack, H. Plank, ACS Omega 2023, 8, 3148 .
- [47] J. D. Fowlkes, R. Winkler, B. B. Lewis, M. G. Stanford, H. Plank, P. D. Rack, ACS Nano 2016, 10, 6163.
- [48] J. Robertson, Mater. Sci. Eng. R 2002, 37, 129.
- [49] L. Keller, M. Huth, Beilstein Journal of Nanotechnology 2018, 9, 2581.
- [50] B. D. Gates, Science 2009, 323, 1566.
- [51] M. Pagliaro, R. Ciriminna, G. Palmisano, Chem. Sus. Chem 2008, 1, 880 .
- [52] K. Rykaczewski, W. B. White, A. G. Fedorov, Journal of Applied Physics 2007, 101, 054307.
- [53] D. C. Joy, Monte Carlo modeling for electron microscopy and microanalysis, volume 9, Oxford University Press 1995.
- [54] R. Browning, T. Z. Li, B. Chui, J. Ye, R. F. W. Pease, Z. Czyżewski, D. C. Joy, J. Appl. Phys. 1994, 76, 2016.
- [55] L. Reimer, Measurement Science and Technology 2000, 11, 1826.
Introduction
Gold is a key metal in microelectronics due to its low resistivity, resistance to electromigration and chemical inertness. These properties make it an excellent choice for stable and reliable electrical contacts even under challenging environmental conditions 1. Chemical vapour deposition (CVD) is a technologically significant option for fabricating gold contacts or other metallic functional structures. CVD utilises the thermal decomposition of precursor molecules in a gaseous state 2. Molecules reaching the surface are thermally decomposed, leaving the central metal atom on the surface. If the decomposition is favored kinetically or thermodynamically on certain surfaces compared to others, the ensuing variation in reaction rates can be exploited for area selective deposition 2. This has been shown for various metals, e.g. Au on W 3, Au on Cr 4, or Cu on Pt 5. All of the aforementioned approaches require sophisticated lithography methods to define the growth surfaces, which can restrict the selection of substrate materials, ultimately constraining the degrees of freedom in device design 3, 6. Consequently, an area selective deposition technique that operates without the need for lithography, and is independent of the substrate type, would be highly advantageous.
One option is to provide a localized energy input for CVD, which is frequently accomplished through laser irradiation 7, 8. However, the spatial resolution achievable through this method is typically restricted by the wavelength of the laser light employed. Although there has been recent evidence that this issue can be addressed in thermally decoupled 3D gold structures 9, the range of achievable structure geometries remains restricted. This is also applicable to local energy input through resistive heating 10. An alternative approach uses beams of charged particles, often electrons 11, 12.
Instead of supplying additional energy during the CVD process, electrons can also be used to activate a specific surface 13, 14 or to deposit a seed layer for preferential decomposition of precursor molecules either by CVD 12, 15 or by atomic layer deposition (ALD) 16, 17. Seeding using a focused electron beam possesses significant potential, as direct electron beam writing is capable of operating on any surface topography that is even weakly conductive 18. Additionally, it provides the possibility for unparalleled spatial resolution 19. If the seeds previously deposited exhibit autocatalytic activity 12, 15, 20, 21, the temperature necessary for decomposition decreases considerably. This technique efficiently limits the CVD metal growth to a specific area, while also reducing the thermal load on the substrate. However, there is currently no proof of electron beam seeding resulting in autocatalytic growth for gold.
In this study, area selective CVD of gold by electron beam seeding is demonstrated for the first time. The process employs a focused electron beam for the non-thermal electron-triggered decomposition of Au(acac)Me which results in autocatalytically active gold nuclei. Subsequently, the same precursor is thermally decomposed at the gold seeds. This autocatalytic CVD of gold is observed at relatively low substrate temperatures of about 80 °C and requires only high vacuum conditions. Electron beam seeding is highly versatile, as demonstrated by the deposition of high aspect ratio pillars using the same precursor. The temperature-dependent autocatalytic CVD in this case provides the first quantitative measurement of the temperature change induced by the electron beam during the direct writing of 3D structures.
Results and Discussion

### v2

@@
Abstract
Chemical vapor deposition (CVD) is an established method for producing high-purity thin films, but it typically necessitates the pre- and post-processing using a mask to produce structures. This paper presents a novel maskless patterning technique that enables area selective CVD of gold. A focused electron beam is used to decompose the metal-organic precursor Au(acac)Me2 locally, thereby creating an autocatalytically active seed layer for subsequent CVD with the same precursor. The procedure could be included in the same CVD process without the need for clean room lithographic processing. Moreover, it operates at low temperatures of 80 °C, over 200 K lower than standard CVD temperatures for this precursor, reducing thermal load on the specimen. Given that electron beam seeding operates on any even moderately conductive surface, the process does not constrain device design. This is demonstrated by the example of vertical nanostructures with high aspect ratios of around 40:1 and more. Written using a focused electron beam and the same precursor, these nanopillars exhibit catalytically active nuclei on their surface. Furthermore, by using the onset of the autocatalytic CVD growth, for the first time the local temperature increase caused by the writing of nanostructures with an electron beam could be precisely determined.
@@ Abstract
area selective chemical vapor deposition, autocatalytic growth, beam-induced heating, direct electron beam writing, goldAleksei Tsarapkin* Krzysztof Maćkosz Chinmai Sai Jureddy Ivo Utke Katja Höflich*
A. Tsarapkin, Dr. K. Höflich
Ferdinand-Braun-Institut (FBH), Leibniz-Institut für Höchstfrequenztechnik, Gustav-Kirchhoff-Str. 4, D-12489 Berlin, Germany
Email Addresses: aleksei.tsarapkin@fbh-berlin.de, katja.hoeflich@fbh-berlin.de
Dr. K. Maćkosz, C. S. Jureddy, Dr. I. Utke
Empa - Swiss Federal Laboratories for Materials Science and Technology, Feuerwerkerstrasse 39, CH-3602 Thun, Switzerland
1 Introduction
@@
Gold is a key metal in microelectronics due to its low resistivity, resistance to electromigration and chemical inertness. These properties make it an excellent choice for stable and reliable electrical contacts even under challenging environmental conditions [1]. Chemical vapor deposition (CVD) is a technologically significant option for fabricating gold contacts or other metallic functional structures. CVD utilises the thermal decomposition of adsorbed precursor molecules delivered in a gaseous state [2]. Molecules reaching the surface are thermally decomposed, leaving the central metal atom on the surface. If the decomposition is favored kinetically or thermodynamically on certain surfaces compared to others, the ensuing variation in reaction rates can be exploited for area selective deposition [2]. This has been shown for various metals, e.g. Au on W [3], Au on Cr [4], or Cu on Pt [5]. All of the aforementioned approaches require sophisticated lithography methods to define the growth surfaces, which can restrict the selection of substrate materials, ultimately constraining the degrees of freedom in device design [3, 6]. Consequently, an area selective deposition technique that operates without the need for lithography, and is independent of the substrate type, would be highly advantageous.
One option is to provide a localized energy input for CVD, which is frequently accomplished through laser irradiation [7, 8]. However, the spatial resolution achievable through this method is typically restricted by the wavelength of the laser light employed. Although there has been recent evidence that this issue can be addressed in thermally decoupled 3D gold structures [9], the range of achievable structure geometries remains restricted. This is also applicable to local energy input through resistive heating [10]. An alternative approach uses beams of charged particles, often electrons [11, 12].
Instead of supplying additional energy during the CVD process, electrons can also be used to activate a specific surface [13, 14] or to deposit a seed layer for preferential decomposition of precursor molecules either by CVD [12, 15] or by atomic layer deposition (ALD) [16, 17]. Seeding using a focused electron beam possesses significant potential, as direct electron beam writing is capable of operating on any surface topography and even on weakly conductive materials [18]. Additionally, it provides the possibility for unparalleled spatial resolution [19]. If the seeds previously deposited exhibit autocatalytic activity [12, 15, 20, 21], the temperature necessary for decomposition decreases considerably. Here, the term autocatalytic CVD is used to highlight that seeds of the same metal enhance the reaction rate for thermal decomposition which efficiently limits the CVD metal growth to a specific area, while also reducing the thermal load on the substrate. However, there is currently no proof of electron beam seeding resulting in autocatalytic growth for gold.
In this study, area selective CVD of gold by electron beam seeding is demonstrated for the first time. The process employs a focused electron beam for the non-thermal electron-triggered decomposition of Au(acac)Me2 which results in autocatalytically active gold nuclei. Subsequently, the same precursor is thermally decomposed at the gold seeds. This autocatalytic CVD of gold is observed at relatively low substrate temperatures of about 80 °C and requires only high vacuum conditions. Electron beam seeding is highly versatile, as demonstrated by the deposition of high aspect ratio pillars using the same precursor. The temperature-dependent autocatalytic CVD in this case provides the first quantitative measurement of the temperature change induced by the electron beam during the direct writing of high aspect-ratio structures.
2 Results and Discussion